OpenStudy (themudkipfryalt):

Identify the equation of the hyperbola with its center at (4, -2), a vertical transverse axis, a = 2, and b = 3.

OpenStudy (amorfide):

\[\frac{ (x-h)^{2} }{ a^{2} } - \frac{ (y-k)^{2} }{ b^{2} } = 1\] where h,k is the centre (h,k) in your case (4,-2)

OpenStudy (amorfide):

http://mathworld.wolfram.com/Hyperbola.html check this out

OpenStudy (mathmale):

amorfide: I'd guess that most students who post a question like this one would need some explanation of what "a" and "b" represent. The Wolfram article, while relevant, is a bit much for a beginner to follow. Secondly, you could help a lot by drawing the basic graphs of vertical and horiz. hyperbolas.

OpenStudy (amorfide):

@mathmale Graphs are not my thing and he had no other help so I thought the equation would be enough Also, he is given the values for a and b, this was the best help I could give, sorry